The question is "Ivan has 80 feet of fence to enclose a rectangular garden. What demensions for the garden give the maximum area?" And I don't know how to solve it so help would be great!
Found 3 solutions by fractalier, Alan3354, ikleyn:
Answer by fractalier(6550)

The answer is a square, 20 x 20 feet.
If you are in calculus, you would take the derivative and set it equal to zero.
If you are in algebra, you could graph it and look for a maximum.
If you are in neither, you could look at how the area changes as you manipulate the dimensions.

Answer by Alan3354(69443)

The question is "Ivan has 80 feet of fence to enclose a rectangular garden. What demensions for the garden give the maximum area?" And I don't know how to solve it so help would be great!
---------------
Max area for a rectangle is a square.
----
P = 2L + 2W = 80
L + W = 40
L = 40 - W
-------------
Area = W*L = W*(40-W)
Area = -W^2 + 40W
The max is the vertex, at W = -b/2a
W = -40/-2 = 20
L = 20

Answer by ikleyn(53765)

.
Let x = the length of the rectangle and y = its width.

Then  x + y = 40  (do you understand why?),  and you are asked to find  x  and  y  in such a way to maximize the product  x*y  which is the area.

Express  y  via  x:  y = 40 - x,  and substitute it into the product:

x*y = x*(40-x)

and find the maximum of the quadratic function  f(x) = x*(40-x).

Can you do it yourself?

If not, please let me know,  I will help you.

Place your answer into the  "Comments from student"  section.  Do not forget to put the number of the problem  (# 1001543)  in order for I could identify it.

Good luck!